OpenStudy (anonymous):

20. Tom, Linda and Alex have $120 dollars. Alex has the third of what Tom has and Linda has twice as much as Alex. How much money, in dollars, does Linda have? A) 10 B) 20 C) 40 D) 60 E) 80

OpenStudy (anonymous):

helo @tester97

OpenStudy (anonymous):

help @tester97

OpenStudy (anonymous):

Equation 1: T + L + A = 120 [Alex Tom and Linda have 120 total] Equation 2: A = 1/3 * T [Alex has 1/3 of what Tom has] This can be re-written as T = 3A Equation 3: L = 2A [Linda has twice as much as Alex] Substitute equations 2 and 3 into equation 1 as follows: T + L + A = 120 [original equation] (3A) + (2A) + A = 120 [substitution] 3A + 2A + A = 6A = 120 so A = 120/6 = 20 Alex has $20, but we're asked to find what Linda has. Plug this answer into Equation 3 to get: L = 2A L = 2(20) L = 40 So Linda has $40

OpenStudy (anonymous):

oh so its C

OpenStudy (tester97):

I concur with @Glorenda49 so yea C

OpenStudy (anonymous):

ok tx

OpenStudy (anonymous):

Glad I could help